### v1

@@ 3.4 Decision-making for Verification
While we establish a method of indicating AEs for a given perturbation $x^{\prime}$ in the last section, to make the decision if the verification target (cf. Def. 3) is met we need to estimate $R_{M}$ over a population of perturbations, for which we propose the “outer loop” in Fig. 3 and explain details in this section.
Concentration inequalities [4], such as Chernoff inequality, Azuma’s bound, and Hoeffding’s inequality, represent important statistical analytical methods extensively employed in ensuring dependable decision-making with probabilistic guarantees. Specifically, in probability theory, Hoeffding’s inequality [19] provides a bound on the probability that the sum of bounded independent random variables deviates from its expected value by more than a certain amount. In ProTIP, for an input $x$, the T2I DM’s probabilistic robustness $R_{M}(x)$ is the proportion of AEs over a population of all possible perturbations on $x$. Normally this population is very large (if not infinite) [^3] The population size depends on the length of the original text $x$ and perturbation parameters like perturbation rate. Note, ProTIP can also cope with extreme (and simpler) cases in which the perturbation population is small and its members can be enumerated, e.g.. when $x$ is a single word with one character to be changed., thus the ground truth of $R_{M}(x)$ can only be estimated as the sample mean of samples drawn from the population. Irrespective of the sample size, there will inevitably be discrepancies between the sample mean and the ground truth population mean. Nevertheless, Hoeffding’s inequality allows us to establish tight probabilistic bounds on this error (cf. Appendix 0.B for more on Hoeffding’s inequality). A limitation of Hoeffding’s inequality is its requirement for a predetermined, process-independent sample size. However, in most circumstances, we generally have no idea how many samples are sufficient to verify the model, a priori. Thus, allocating a large sample size that accommodates the budget can result in unnecessary waste.
@@
The proof for Theorem 3.1 is presented in [62], which adapts the more general proof in [63] for the binary 0–1 variables $I_{i}$ and also rearranges terms to align with the form of the original Hoeffding’s inequality. We further prove two corollaries in Appendix 0.B to better explain our later experimental results: One compares the tightness of the two bounds derived from the original and adaptive Hoeffding’s inequality; the other concerns their monotonicity with respect to $n$ and $\sigma$.
In ProTIP, we directly apply the result of Eq. (6) by sequentially increasing the number of perturbations from $J=1$ to $J=j_{max}$ (the maximum sampling budget for generating perturbations). Given the verification target $(b_{l},\sigma)$ (cf. Def. 3), whenever the empirically estimated lower bound on $R_{M}(x)$ yielded by Eq. (6) (after rearranging the inequality) is greater or equal to the specified requirement $b_{l}$ (i.e., $\hat{\mu}_{I}^{(J)}-\varepsilon(\sigma,J)\geq b_{l}$), we stop generating new perturbations and assert “pass” for the verification. Otherwise, when $J=j_{max}$, ProTIP asserts “fail” with an evaluated robustness.
@@ 4 Experiments
To evaluate the effectiveness and efficiency of the proposed framework, ProTIP, we conduct experiments with the following setup.
We choose two versions of the widely acclaimed and open-source [^4] We conduct experiments locally on our own servers for efficiency, as ProTIP generates a large number of queries to the model under verification. Therefore, we exclude experiments on non-open source models, such as DALLE-3 and Midjourney. Stable Diffusion (SD) model [41], SD V1.5 and SD V1.4. While different versions of SD models have the same structure, the higher version is further trained based on the previous version. We use the MS-COCO dataset [28], a comprehensive collection designed for common computer vision tasks including captioning. The dataset comprises 328,000 images with captions, from which we randomly select image captions as prompts for the T2I DM and then apply the stochastic perturbation method discussed in Sec. 3.2 on these prompts to generate the perturbed inputs.
All the experiments are conducted on NVIDIA GeForce RTX 3090 GPUs, Python 3.11, PyTorch 2.1.1. We use the off-the-shelf R package [39] to design group sequential analysis parameters, cf. Appendix 0.C for details. All our models, data, source code and experimental results are publicly accessible at https://github.com/wellzline/ProTIP/.
@@ 4.1 Effectiveness of ProTIP
In order to evaluate the effectiveness of ProTIP, i.e., how accurate our ProTIP is, we need to know the ground truth probabilistic robustness $R_{M}(x)$ for the given prompt $x$. However, as articulated in Sec. 3.1, the ground truth of $R_{M}(x)$ can never be known (as it would require exhaustively testing all possible perturbations). Consequently, we can only approximate the ground truth by using a significantly larger number of samples than what would normally be used in ProTIP to demonstrate accuracy, which is arguably a common practice in statistical inference, e.g., [53, 20, 8]. Specifically, we conduct the following two steps: i) estimate the ground truth of $I(x^{\prime})$ for a given perturbation $x^{\prime}$ by generating a large number of images; ii) then estimate the ground truth $R_{M}(x)$ using a large number of $x^{\prime}$s with their ground truth $I(x^{\prime})$. For step 1, we stop generating the images for the $x^{\prime}$ until the distribution of its CLIP scores “converge” (i.e., the shape of the CLIP score distribution shows no significant changes when new images are generated). We then check whether the CLIP scores, as well as the CLIP score distribution of the original prompt $x$, follow a normal distribution. If they do, we execute a t-test, otherwise a u-test. In step 2, we generate a fixed number of $1,000$ perturbations, which is much larger than the adaptive number of perturbations (around $50\!\sim\!350$) used in ProTIP. Then the original Hoffeding’s inequality, which yields a tighter error bound (cf. Corollary 1) than the adaptive version in ProTIP, is applied to approximate the ground truth $R_{M}(x)$.
@@
Fig. 4(b) presents the results with an increased perturbation rate of 20%. As expected, the ground truth robustness is lower, as well as the ProTIP results—it asserts “unverified” reflects that the ground truth is lower than the verification target. In Fig. 4(c), ProTIP is applied to a older version SD V1.4 model. Compared to (a), it is evident that, under the same 10% perturbation rate, the estimated robustness of SD V1.4 are lower than those of SD V1.5, implying its weaker ability to withstand adversarial perturbations, which conforms to the observation in [61, 26, 49]. While Fig. 4 only shows the results of a single prompt, results for more prompts are presented in Appendix 0.D.
Figure 4: ProTIP results for a given prompt, with different confidence levels $1-\sigma$.
@@ 4.2 Efficiency of ProTIP
The efficiency of ProTIP manifests in two aspects, firstly in identifying AEs, as illustrated in the “inner loop” of Fig. 3. The aforementioned challenge of identifying a perturbed input as an AE involves comparing distributional differences which is computationally insensitive for T2I models. Thus, we employ sequential analysis with efficacy and futility early stopping rules in the statistical testing for indicating AEs. Following the parameter settings of Appendix 0.C, WLOG, we divide the sequential analysis into 5 stages to conduct interim analysis.
@@
Fig. 4(a) illustrates the number of perturbations indicated as Non-AEs/AEs at each interim analysis stage, for a given prompt processed by SD V1.5 with perturbation rates of 10% and 20%, as well as by SD V1.4 with a 10% perturbation rate. For V1.5 with 10% rate, we observe that, out of the total 400 perturbations, 347 are identified as Non-AEs. Thanks to the futility stopping rule, 246 of these Non-AEs can be identified at stage 1, saving 4 times the cost of generating images compared to what would be required without early stopping rules (i.e., test at the final stage 5). Considering other stages as well, in total it reduces the computational cost by detecting 97% of Non-AEs before reaching the final stage 5. Similarly, 68% of AEs are detected early due to the efficacy stopping rule.
With a 20% perturbation rate and an older version, the total number of AEs increased and Non-AEs decreased, yet overall, the effect of two early stopping rules in the statistical testing for indicating Non-AEs/AEs remains evident. While Fig. 4(a) demonstrates the results for only one prompt, statistics for 36 randomly selected prompts are presented in Table 2.
@@
The other efficient aspect of ProTIP is demonstrated by introducing the adaptive concentration inequalities to dynamically determine the “just-right” number of perturbations whenever the verification target is met, cf. the “outer loop” of Fig. 3. To compare with the original Hoeffding’s inequality which requires a predetermined, process-independent sample size, we have to do “what if” calculations by assuming the sample size used by it. Fig. 4(b) presents the ProTIP results and the (original) Hoeffding’s inequality with sample sizes of 50, 200, and 400 (vertical dashed lines in green). With sample size 50, Hoeffding’s inequality asserts an incorrect verification result “fail”, while ProTIP correctly verifies it, although with more samples of 225. This is non-surprising as the error bound is bigger when with limited samples (cf. Corollary 2). Thus no practitioners would apply Hoeffding’s inequality with such small sample size, rather allocate a much larger sample size for a small estimation error. The case with sample size 400 is replicating this scenario, in which we observe that both methods make the correct verification decision while ProTIP saves 175 perturbations. For the case of sample size 200, indeed both methods yield similar results (and Hoeffding’s inequality is slightly better, cf. Corollary 1) using a similar number of samples. However, in practice, we would never know the “just right” sample size a priori when applying Hoeffding’s inequality, highlighting ProTIP’s superiority in terms of efficiency.
@@
(a) Number of perturbations (out of 400 perturbations) identified as Non-AEs/AEs at each interim stage (S) in the sequential hypothesis testing.
(b) ProTIP with the (adaptive) sample size of 225 vs. Hoeffding’s inequality with fixed sample size 50, 200, and 400.
Figure 5: Illustration of ProTIP efficiency for a given prompt.
@@
Table 2: The mean percentage of perturbations indicated as Non-AE/AE by the two early stopping rules before the final stage 5, over 36 randomly selected prompts.
@@ 4.3 Application of ProTIP for Ranking Defence Methods
As a robustness assessment tool, a natural use case of our ProTIP is to rank common defence methods for T2I DMs. For character-level perturbations that are perceivable and semantic, scrutinising the input is a direct and universally applicable passive defence. Thus, we study three commonly used misspelling checking tools: the Python Autocorrect 0.3.0 package [12], the Pyspellchecker package [36], and Gramformer [37] for grammar error correction (GEC) on the perturbed inputs before inputting them into the T2I DM.
Figure 6: Box plot of the probabilistic robustness with/without defence methods, over 36 randomly selected prompts.
In Fig. 7, for an example prompt, all three defence tools may improve the model’s resistance to stochastic perturbations in input text, compared to the the case without using any defenders (the red curve). Notably, Autocorrect demonstrates superior overall performance compared to the other two tools. Such observations are further confirmed by the box plots in Fig. 6 over a set of 36 randomly selected prompts. Our approach allows for the performance ranking of these misspelling checking tools, as shown in the Fig. 6.
@@ 5 Conclusion
In this work, for the first time, we formalise the definition of probabilistic robustness of T2I DMs and then establish an efficient framework, ProTIP, for evaluating it. As a black-box verification method, ProTIP is based on principled statistical inference approaches. It incorporates sequential analysis with efficacy and futility early stopping rules in the statistical hypothesis testing for identifying AEs, along with adaptive concentration inequalities to dynamically adjust the number of stochastic perturbations needed to make the verification decision. Empirical experiments have substantiated the effectiveness and efficiency of ProTIP. Finally, we demonstrate a use case of ProTIP to rank various commonly used defence methods, highlighting its versatility and applicability.
Appendix 0.A Text-to-Image Diffusion Model
Diffusion models are probabilistic generative models that are trained to learn a data distribution by the gradual denoising of a variable sampled from a Gaussian distribution. One specific latent diffusion architecture is shown in Fig. 8. Specifically, for a pre-trained T2I DM $\hat{x}_{\theta}$, given an initial noise map $\gamma\sim\mathcal{N}(0,I)$ and a conditioning vector $c=\Gamma(P)$ generated using a text encoder $\Gamma$ and a text input $P$, generates an image $x_{\text{gen}}=\hat{x}_{\theta}(\gamma,c)$. For text-to-image generation, the model is conditioned on the text prompt $P$ and guided toward an image faithful to that prompt. The training objective of a diffusion model $\hat{x}_{\theta}$ can be written as:
$E_{x,c,\gamma,t}[w_{t}||\hat{x}_{\theta}](\alpha_{t}x+\sigma_{t}\boldsymbol{\gamma},c)-x||^{2}_{2}]$ (7)
Where $x$ is the ground-truth image, $c$ is a conditioning vector (e.g., obtained from a text prompt), and $\alpha_{t}$, $\sigma_{t}$, $w_{t}$ are terms that control the noise schedule and sample quality, and are functions of the diffusion process time $t\sim\mathcal{U}([0,1])$. Consequently, the DM is trained to denoise $z_{t}:=\alpha_{t}x+\sigma_{t}\gamma$ yielding x with the squared error loss. At inference time, the diffusion model is sampled by iteratively denoising $z_{t_{1}}\sim\mathcal{N}(0,I)$ using either the deterministic DDIM [44] or the stochastic ancestral sampler [18]. Intermediate points $z_{t_{1}},\ldots,z_{t_{T}}$, where $01=t_{1}>\ldots>t_{T}=$ 0, are generated, with decreasing noise levels. These points, $\hat{x}_{0}^{(t)}:=\hat{x}_{\theta}(z_{t},c)$, are functions of the $x$-predictions.
Figure 8: Overview of the latent diffusion architecture. Image from the paper[41].
Appendix 0.B Hoeffding’s Inequality

### v2

@@ 3.4 Decision-making for Verification
While we establish a method of indicating AEs for a given perturbation $x^{\prime}$ in the last section, to make the decision if a given verification target (cf. Def. 3) is met we need to estimate $R_{M}$ over a population of perturbations, for which we propose the “outer loop” in Fig. 3 and explain details in this section.
Concentration inequalities [4], such as Chernoff inequality, Azuma’s bound, and Hoeffding’s inequality, represent important statistical methods widely employed in ensuring dependable decision-making with probabilistic guarantees. Specifically, in probability theory, Hoeffding’s inequality [20] provides a bound on the probability that the sum of bounded independent random variables deviates from its expected value by more than a certain amount. In ProTIP, for an input $x$, the T2I DM’s probabilistic robustness $R_{M}(x)$ is the proportion of AEs over a population of all possible perturbations on $x$. Normally this population is very large (if not infinite) [^3] The population size depends on the length of the original text $x$ and perturbation parameters like perturbation rate. Note, ProTIP can also cope with extreme (and simpler) cases in which the perturbation population is small and its members can be enumerated, e.g., when $x$ is a single word with one character to be changed., thus the ground truth of $R_{M}(x)$ can only be estimated as the sample mean of samples drawn from the population. Irrespective of the sample size, there will inevitably be discrepancies between the sample mean and the ground truth population mean. Nevertheless, Hoeffding’s inequality allows us to establish tight probabilistic bounds on this error (cf. Appendix C for more on Hoeffding’s inequality). A limitation of Hoeffding’s inequality is its requirement for a predetermined, process-independent sample size. However, in most circumstances, we generally have no idea how many samples are sufficient to verify the model, a priori. Thus, it is common to allocate a maximised sample size that accommodates the budget limit which may result in unnecessary waste.
@@
The proof for Theorem 3.1 is presented in [66], which adapts the more general proof in [67] for the binary 0–1 variables $I_{i}$ and also rearranges terms to align with the form of the original Hoeffding’s inequality. We further prove two corollaries in Appendix C to better explain our later experimental results: One compares the tightness of the two bounds derived from the original and adaptive Hoeffding’s inequality; the other concerns their monotonicity with respect to $n$ and $\sigma$.
In ProTIP, we directly apply Eq. (6) by sequentially increasing the perturbation number $J$ from 1 to $j_{max}$ (the maximum sampling budget). Given the verification target $(b_{l},\sigma)$ (cf. Def. 3), whenever the empirically estimated lower bound on $R_{M}(x)$ yielded by Eq. (6) (after rearranging the inequality) is greater or equal to the specified requirement $b_{l}$ (i.e., $\hat{\mu}_{I}^{(J)}-\varepsilon(\sigma,J)\geq b_{l}$), we stop generating new perturbations and assert “pass” for the verification. Otherwise, when $J=j_{max}$, ProTIP asserts “fail” with an estimated robustness.
@@ 4 Experiments
In our experiments, we study three versions of the widely acclaimed and open-source [^4] We run experiments locally on our own servers for efficiency, as ProTIP generates a large number of queries to the model under verification. Thus, we exclude experiments on non-open source, commercial models, e.g., DALLE-3 and Midjourney. Stable Diffusion (SD) model [42], SD-V1.5, SD-V1.4 and SDXL-Turbo. While different versions of SD models have the same structure, the higher version is further trained based on the previous version; and SDXL-Turbo is based on Adversarial Diffusion Distillation [45]. We use the MS-COCO dataset [30], a comprehensive collection designed for common computer vision tasks including captioning. The dataset comprises 328,000 images with captions, from which we randomly select captions as prompts for the T2I DM and then apply the stochastic perturbation method discussed in Sec. 3.2 to generate perturbed inputs.
All experiments are conducted on NVIDIA GeForce RTX 3090 GPUs, Python 3.11, PyTorch 2.1.1. We use the off-the-shelf R package [40] to design group sequential analysis parameters, cf. Appendix D for details. All our models, data, source code and experimental results are publicly accessible at our Github site.
@@ 4.1 Effectiveness of ProTIP
In order to evaluate the effectiveness of ProTIP, i.e., how accurate our ProTIP is, we need to know the ground truth probabilistic robustness $R_{M}(x)$ for the given prompt $x$. However, as articulated in Sec. 3.1, the ground truth of $R_{M}(x)$ can never be known (as it would require exhaustively testing all possible perturbations). Consequently, we can only approximate the ground truth by using a significantly larger number of samples than what would normally be used in ProTIP to demonstrate accuracy, which is arguably a common practice in statistical inference, e.g., [56, 21, 8]. Specifically, we conduct the following two steps: i) estimate the ground truth of $I(x^{\prime})$ for a given perturbation $x^{\prime}$ by generating a large number of images; ii) then estimate the ground truth $R_{M}(x)$ using a large number of $x^{\prime}$s with their ground truth $I(x^{\prime})$. For step 1, we stop generating the images for the $x^{\prime}$ until the distribution of its CLIP scores “converge” (i.e., the shape of the CLIP score distribution shows no significant changes when new images are generated). We then check whether the two CLIP score distributions follow a normal distribution. If yes, we execute a t-test, otherwise a u-test. In step 2, we generate a fixed number of $1,000$ perturbations, which is much larger than the adaptive number of perturbations (around $50\!\sim\!350$) used in ProTIP. Then the original Hoffeding’s inequality (which yields a tighter error bound than the adaptive version used in ProTIP for the same number of perturbations, cf. Corollary 1) is applied to approximate the ground truth $R_{M}(x)$.
@@
Fig. 4(b) presents the results with an increased perturbation rate of 20%. As expected, the ground truth robustness is lower, as well as the ProTIP results—it asserts “fail” reflects that the ground truth is lower than the verification target. In Fig. 4(c), ProTIP is applied to a older version SD V1.4 model. Compared to (a), it is evident that, under the same 10% perturbation rate, the estimated robustness of SD V1.4 are lower than those of SD V1.5, implying its weaker ability to withstand adversarial perturbations, which conforms to the observation in [65, 28, 52]. While Fig. 4 only shows the results of a single prompt, results for more prompts and SDXL-Turbo are presented in Appendix E
Figure 4: ProTIP results for a given prompt, with different confidence levels $1-\sigma$ (a)(b); Mean & Std. (shared area) of RLB over 107 prompts (c).
@@ 4.2 Efficiency of ProTIP
The efficiency of ProTIP manifests firstly in identifying AEs—the “inner loop” of Fig. 3. As the aforementioned sequential statistical testing method, following the parameter settings of Appendix D, WLOG we divide it into 5 stages to conduct interim analysis. Fig. 5(a) illustrates the number of perturbations identified as Non-AEs/AEs at each interim analysis stage, for a given prompt processed by SD V1.5 with perturbation rates of 10% and 20%, as well as by SD V1.4 with a 10% perturbation rate. For V1.5 with 10% rate, we observe that, out of the total 400 perturbations, 347 are identified as Non-AEs. Thanks to the futility stopping rule, 246 of these Non-AEs can be identified at stage 1, saving 4 times the cost of generating images compared to what would be required without early stopping rules (i.e., test at the final stage 5). Considering other stages as well, in total it reduces the computational cost by detecting 97% of Non-AEs before reaching the final stage 5. Similarly, 68% of AEs are detected early due to the efficacy stopping rule. With a 20% perturbation rate and an older version, the total number of AEs increased and Non-AEs decreased, yet overall, the effect of two early stopping rules in the statistical testing for indicating Non-AEs/AEs remains evident. While Fig. 5(a) demonstrates the results for only one prompt, statistics for 36 randomly selected prompts are presented in Table 1.
@@
The other efficient aspect of ProTIP is demonstrated by introducing the adaptive concentration inequalities to dynamically determine the number of perturbations, cf. the “outer loop” of Fig. 3. To compare with the original Hoeffding’s inequality which requires a predetermined, process-independent sample size, we have to do “what if” calculations by assuming the sample size used by it. Fig. 5(b) presents the ProTIP results and the (original) Hoeffding’s inequality with sample sizes of 50, 100, and 200 (dotted cross in green). With sample size 50, Hoeffding’s inequality asserts an incorrect verification result “fail”, while ProTIP correctly verifies it, although with more samples of 78. This is non-surprising as the error bound is bigger when with limited samples (cf. Corollary 2). Thus no practitioners would apply Hoeffding’s inequality with such small sample size, rather allocate a much larger sample size for a small estimation error. The case with sample size 200 is replicating this scenario, in which both methods make the correct verification decision while ProTIP saves 122 perturbations. For the case of sample size 100, indeed both methods yield similar results (and Hoeffding’s inequality is slightly better, cf. Corollary 1) with similar numbers of samples. However, in practice, we never know such “just right” sample size a priori when applying Hoeffding’s inequality, highlighting ProTIP’s superiority in efficiency.
@@
Figure 5: (a) Number of perturbations (out of 400) identified as Non-AEs/AEs at each interim stage (S) in the sequential hypothesis testing. (b) ProTIP with the (adaptive) sample size of 78 vs. Hoeffding’s inequality with fixed sample size 50, 100, and 200.
@@
Table 1: The mean and std. of perturbations identified as Non-AE/AE by the two early stopping rules before the final stage 5, over 36 randomly selected prompts.
@@ 4.3 Application of ProTIP for Ranking Defence Methods
As a robustness assessment tool, a natural use case of our ProTIP is to rank common defence methods for T2I DMs. For character-level perturbations that are perceivable and semantic, scrutinising the input is a direct and universally applicable defence. Thus, we study three commonly used misspelling checking tools: Python Autocorrect 0.3.0 [12], Pyspellchecker [37], and Gramformer [38] for error correction on the perturbed inputs before inputting them into T2I DMs.
In Fig. 6, for an example prompt, all three defence tools may improve the model’s resistance to perturbations in inputs, compared to the case without using any defenders (the red curve). Notably, Autocorrect demonstrates superior overall performance compared to the other two tools. Such observations are further confirmed by the box plots in Fig. 7 over a set of 36 randomly selected prompts, ranking their performance. Cf. Appendix E for results of more prompts/models.
@@ 5 Conclusion
Figure 7: Box plot of the probabilistic robustness with & without defenders, over 36 randomly selected prompts.
@@
In this work, for the first time, we formalise the definition of probabilistic robustness for T2I DMs and then establish an efficient framework, ProTIP, for evaluating it. As a black-box verification method, ProTIP is based on principled statistical inference approaches. It incorporates sequential analysis with early stopping rules in hypothesis testing when identifying AEs, along with adaptive concentration inequalities to dynamically adjust the number of stochastic perturbations needed to make the verification decision. Empirical experiments have substantiated the effectiveness and efficiency of ProTIP. Finally, we demonstrate a use case of ProTIP to rank various commonly used defence methods, highlighting its versatility and applicability.
Acknowledgments
This work is supported by the British Academy’s Pump Priming Collaboration between UK and EU partners 2024 Programme [PPHE24/100115]. We thank Hongchen Gao for his valuable comments on our work.
Appendix 0.A Denoising Diffusion Probabilistic Models
Denoising Diffusion Probabilistic Models (DDPMs)[19] [25] [46] are a class of probabilistic generative models that apply a noise injection process, followed by a reverse procedure for sample generation. A DDPM is defined as two parameterized Markov chains: a forward chain that add random Gaussian noise to images to transform data distribution into a simple prior distribution and a reverse chain that convert the noised image back into target data by learning transition kernels parameterized by deep neural networks.
Forward diffusion process: Given a data point sampled from a real data distribution $x_{0}\sim q(x)$, a forward process begins with adding a small amount of Gaussian noise to the sample in $T$ steps, producing a sequence of noisy samples $x_{1},\ldots,x_{T}$. The step sizes are controlled by a variance schedule $\{\beta_{t}\in(0,1)\}_{t=1}^{T}$.
$q(x_{1:T}|x_{0})$ $:=\prod_{t=1}^{T}q(x_{t}|x_{t-1}),$ (7)
$q(x_{t}|x_{t-1})$ $=\mathcal{N}\left(x_{t};(1-\beta_{t})x_{t-1},\beta_{t}I\right).$
The data sample $x_{0}$ gradually loses its distinguishable features as the step $t$ becomes larger. Eventually, when $T\to\infty$, $x_{T}$ is equivalent to an isotropic Gaussian distribution.
Reverse diffusion process: The reverse process starts by first generating an unstructured noise vector from the prior distribution, then gradually removing noise by running a learnable Markov chain in the reverse time direction. Specifically, the reverse Markov chain is parameterized by a prior distribution $p(x_{T})=\mathcal{N}(x_{T};0,I)$ and a learnable transition kernel $p_{\theta}(x_{t-1}|x_{t})$. Therefore, we need to learn a model $p_{\theta}$ to approximate these conditional probabilities in order to run the reverse diffusion process.
$p_{\theta}(x_{0:T})$ $=p(x_{T})\prod_{t=1}^{T}p_{\theta}(x_{t-1}|x_{t}),$ (8)
$p_{\theta}(x_{t-1}|x_{t})$ $=\mathcal{N}\left(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t)\right),$
where $\theta$ denotes model parameters, often instantiated by architectures like UNet, which parameterize the mean $\mu_{\theta}(x_{t},t)$ and variance $\Sigma_{\theta}(x_{t},t)$. The UNet takes the noised data $x_{t}$ and time step $t$ as inputs and outputs the parameters of the normal distribution, thereby predicting the noise $\epsilon_{\theta}$ that the model needs to reverse the diffusion process. With this reverse Markov chain, we can generate a data sample $x_{0}$ by first sampling a noise vector $x_{T}\sim p(x_{T})$, then successively sampling from the learnable transition kernel $x_{t-1}\sim p_{\theta}(x_{t-1}|x_{t})$ until $t=1$.
Appendix 0.B Stochastic Perturbation Generation
CLIP Score We summarise the following reasons for choosing CLIP Score to measure the correlation between a generated caption for an image and the actual content of the image.
  • 1. While CLIP score is certainly not perfect as a metric (like other metrics) to mimic human-perception similarity, but some works show that CLIP is generally reliable and highly correlated with human judgement in user experiments, e.g.,[18]. Both CLIP scores are calculated (and thus anchored) by considering the text information, regardless of variations in image styles and backgrounds for additional evidence. Therefore, they are highly likely to be similar.
  • 2. While the best metric for measuring T2I correctness remains an open question, CLIP score is commonly used in studying T2I robustness [11, 69, 64, 31, 47, 63]
  • 3. While CLIP score is a building block of ProTIP, it can be substituted with other correctness metrics. The statistical models—the theoretical core and main contribution of ProTIP—are easily adaptable to any new correctness metrics as future advancements.
Stochastic Perturbation Method Table 2 for examples of stochastic text perturbations on T2I DM inputs.
Table 2: Examples of stochastic text perturbations
Perturbation Description Example
Insert Insert a character randomly A white daog plays with a red ball on the green grass.
Substitute Substitute a character randomly A white dog plays with a rad ball on the green grass.
Swap Swap two characters randomly A white dog plays with a red ball on the green garss.
Delete Delete a character randomly A white dog plays with a red bll on the green grass.
Keyboard Substitute a char. by keyboard distance A whote dog plays with a red ball on the green grass.
Appendix 0.C Hoeffding’s Inequality